OpenStudy (anonymous):

Use the quadratic formula to find any x-intercepts on the graph of the equation. y=x^2+6x-10

OpenStudy (anonymous):

equate with 0

OpenStudy (anonymous):

x^2+6x-10=0

OpenStudy (anonymous):

find values of x, and that's the intersections

OpenStudy (jack1):

-b+or-sqrt ((b^2-4ac)/2a)

OpenStudy (anonymous):

okay, after setting it up to = 0 you have to factor 6 and 10 correct?

OpenStudy (anonymous):

it's all up to you, whatever method you choose

OpenStudy (anonymous):

yes, you can do it like that

OpenStudy (anonymous):

okay well they have common factors of 2..but I dont know what makes 6 but multiples into 10?

OpenStudy (jack1):

\[y = a x ^{2} + bx + c\]

OpenStudy (anonymous):

yeah, this method can not move you further, use \[x=\frac{ -b+\sqrt{b ^{2}-4ac} }{ 2 }\]

OpenStudy (anonymous):

Jack1, have shown

OpenStudy (jack1):

dude its over 2a, not 2

OpenStudy (jack1):

and its plus or minus

OpenStudy (anonymous):

yep, it's so buggy

OpenStudy (jack1):

cool website though

OpenStudy (jack1):

peckce how're u doing?

OpenStudy (anonymous):

Sorry, I am just trying to put all the info in the formula..I really hate this formula.

OpenStudy (anonymous):

@Jack1, yes website is great, but it has leaks

OpenStudy (anonymous):

sometimes i can not enter it, xD

OpenStudy (anonymous):

okay so it should look like this, \[-6\pm \sqrt{6} ^{2}-4(1)(-10) / 2(1)\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

leaving us with \[-6\pm \sqrt{36-40}/ 2 = -6\pm \sqrt{-4}/2\]

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

36+40

OpenStudy (anonymous):

so 76

OpenStudy (anonymous):

\[\sqrt{76}= 39?\]

OpenStudy (anonymous):

no and no

OpenStudy (anonymous):

dang.

OpenStudy (anonymous):

−3±√14

OpenStudy (anonymous):

could you show me how that works?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

how works formula u mean?

OpenStudy (anonymous):

no how you got\[-3\pm \sqrt{14} \]

OpenStudy (anonymous):

do you see the formula above?

OpenStudy (anonymous):

Jack1 told that ax^2+bx+c=0, then mine

OpenStudy (anonymous):

yeah i seen it and last I had was \[-6\pm \sqrt{76}/2 \]

OpenStudy (anonymous):

\[\sqrt{76}=\sqrt{4*14}=2\sqrt{14}\]

OpenStudy (anonymous):

so...

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yes and then we divide 2 into the \[-6 \pm \]?

OpenStudy (anonymous):

yes

OpenStudy (jack1):

dont you mean sqrt 19...?

OpenStudy (anonymous):

therefore it leaves us with \[-3\pm 1\sqrt{14} ??\]

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

How is the Sqrt of 19?

OpenStudy (jack1):

my intercepts were approx 1.36 and -7.36 ish exact answers ((sqrt19) - 3) and -((sqrt19) +3) is that what you got pecke?

OpenStudy (anonymous):

I was at sqrt of 76/2 since then i've been lost.

OpenStudy (jack1):

as sqrt 76 is 2 x sqrt 19

OpenStudy (anonymous):

click on my atachment for the answer

OpenStudy (anonymous):

@cmm I did, the graph just should the point at -9... I am unsure of how it's answer when I need it in the radical form.

OpenStudy (anonymous):

therefore would the new answer be \[-3\pm 1 \sqrt{19}\]

OpenStudy (anonymous):

x = 3

OpenStudy (jack1):

[-6 +/- (sqrt 76)] / 2 =-3 +/- (sqrt 76)/2 =-3 +/- (2x(sqrt 19))/2 =-3 +/- (sqrt 19) 2 x intercepts so 2 answers that's where the plus OR minus comes in, one for each intercept

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but how come it doesn't have it showing 2/2 = 1 giving it -3+/- 1 sqrt 19?

OpenStudy (jack1):

you can have 1x sqrt 19... but it still just equals sqrt 19...

OpenStudy (anonymous):

sorry i miscalculated((

OpenStudy (anonymous):

Okay, I only ask because my last question like this one, showed the 1

OpenStudy (anonymous):

Nevermind I was mistaken it didn't

OpenStudy (jack1):

i think it's a waste of space writing the one and it may be mistaken for subscript on the root factor (ie cube root is \[\sqrt[3]{?}\]

OpenStudy (anonymous):

their are two answers (0,0) and (6,0) as the x intercept

OpenStudy (anonymous):

okay so final answer should look like this. \[-3+\sqrt{19}, -3-\sqrt{19}\]

OpenStudy (jack1):

yep, perfect if you want exact answers, 1.36 and -7.36 if you're just looking to plot it by hand

OpenStudy (anonymous):

nope, thankfully don't have to do that. thank you guys!

OpenStudy (jack1):

sall good, happy dayz